OpenStudy (anonymous):

I need help !!); x^3/5 ------- x^1/4

OpenStudy (anonymous):

Divide

OpenStudy (owlcoffee):

\[\frac{ \frac{ x^3 }{ 5 } }{ \frac{ x^1 }{ 4 } }\] Whenever you deal with a division of fractions, the technique to divide them is quite simple, you flip one of them and multiply them instead: \[\frac{ \frac{ a }{ b } }{ \frac{ x }{ y } }=(\frac{ a }{ b })(\frac{ y }{ x })\] So, applying it: \[\frac{ \frac{ x^3 }{ 5 } }{ \frac{ x^1 }{ 4 } } =(\frac{ x^3 }{ 5 })(\frac{ 4 }{ x^1 })\]

OpenStudy (anonymous):

7/20?

OpenStudy (owlcoffee):

That's incorrect, "x" is a variable and you should multiply the constant with it. Notice how there is another x, on the denominator you can cancel: \[(\frac{ x^3 }{ 5 })(\frac{ 4 }{ x^1 })=(\frac{ x^2 }{ 5 })(\frac{ 4 }{ 1 })\]

OpenStudy (anonymous):

Oh got it thank you!